OpenStudy (anonymous):

How would I do this???? Help? The length of a rectangle is 1ft more than twice its width, and the area of the rectangle is 28ft^2 . Find the dimensions of the rectangle.

OpenStudy (ash2326):

Let the width of the rectangle be x We have the information: "The length of a rectangle is 1ft more than twice its width," Then the length would be 2x+1 Can you try to solve for x?

OpenStudy (anonymous):

3x

OpenStudy (anonymous):

width=x and length=2x+1 area of rectangle=lenth*width x(2x+1)=28 2x^2+x-28=0 or x^2+1-14=0

OpenStudy (anonymous):

sorry x^2+x-14=0

OpenStudy (anonymous):

now solve it to get the x

OpenStudy (anonymous):

sorry the equation should be 2x^2+x-28=0

OpenStudy (anonymous):

width =7 and length=15

OpenStudy (anonymous):

sorry width=3.5 and length= 8